OpenStudy (anonymous):

Find f if grad f = (1/x)+(1/y)+(1/z).

OpenStudy (accessdenied):

Is it more like: \( \displaystyle \nabla f = \frac{1}{x} \hat{i} + \frac{1}{y} \hat{j} + \frac{1}{z} \hat{k} \) With those components? I think if that's how it was written, the definition of gradient could be used with equating the like-components. \( \displaystyle\nabla f = \frac{\partial f}{\partial x} \hat{i} + \frac{\partial f}{\partial y} \hat{j} + \frac{\partial f}{\partial z} \hat{k} \)

OpenStudy (anonymous):

yup, which looks like it should be ln(x)+ln(y)+ln(z)+c, but webassign begs to differ...

OpenStudy (accessdenied):

well, the partial derivatives seem to bring up an interesting case, where you'd actually lose little function pieces of the other variables rather than a constant in partial differentiation... Like if f is a function of x, y, and z, you'd lose any of the pieces without an x if you take that partial derivative w.r.t. x.

OpenStudy (accessdenied):

...Although, I am having trouble thinking of what makes your answer incorrect, since that solution does seem to work out, and I can't think of any way to find another working solution. I am wondering if what you're answering this in has a specific handling on the constant you add in, if that's the standard indication of an additive constant or not. That could 'potentially' be a reason...

OpenStudy (accessdenied):

\[ \frac{\partial f}{\partial x} = \frac{1}{x} \\ \frac{\partial f}{\partial y} = \frac{1}{y} \\ \frac{ \partial f}{\partial z} = \frac{1}{z} \\ ~\\ f = \ln x + g(y, z) \\ \frac{\partial f}{\partial y} = \frac{1}{y} \\ \frac{\partial f}{\partial z} = \frac{1}{z} \\ ~\\ \frac{\partial}{\partial y} \left( \ln x + g(y, z) \right) = \frac{1}{y} \\ \frac{\partial}{\partial z} \left( \ln x + g(y, z) \right) = \frac{1}{z} \\ ~\\ \frac{\partial g}{\partial y} = \frac{1}{y} \\ \frac{\partial g}{\partial z} = \frac{1}{z} \\ ~\\ g = \ln y + h(z) \\ \frac{\partial}{\partial z} \left(\ln y + h(z) \right) = \frac{1}{z} \\ \frac{\partial h}{\partial z} = \frac{1}{z} \\ h = \ln z + c \\ g = \ln y + \ln z + c \\ f = \ln x + \ln y + \ln z + c \] Yeah, I get the same answer here...